Circular No. 5946

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994C IN ANONYMOUS GALAXY J. Mueller reports her discovery of a supernova of mag 17.5-18, found on

a blue plate taken on Mar. 5 UT with the 1.2-m Oschin Schmidt Telescope by J. D. Mendenhall and herself in the course of the second Palomar Sky Survey. SN 1994C is located at α = 7^{h}53^{m}08^{s}.07, δ = +45^{o}00′20′′.8 (equinox 1950.0); the galaxy’s center is located at (end figures) 08^{s}.27, 14′′.1. A spectrogram obtained by S. Djorgovski, D. Thompson, and J. Smith on Mar. 9 with the 5-m Hale Telescope (+ double spectrograph) confirms that the object is a supernova.

SUPERNOVA 1994D IN NGC 4526 R. R. Treffers, A. V. Filippenko, and S. D. Van Dyk, University of

California at Berkeley; and M. W. Richmond, Princeton University, report their discovery of a supernova in NGC 4526 (α = 12^{h}31^{m}31^{s}, δ = +7^{o}58′.5, equinox 1950.0), located ∼ 9′′ west and 7′′ north of the galaxy’s nucleus: “The object was found on Mar. 7 UT (R = 15.2 ± 0.5) during the Leuschner Observatory Supernova Search, which uses an automated 0.76-m telescope equipped with the Lawrence Berkeley Laboratory CCD camera. Confirmation was made with an image taken through thin fog on Mar. 9, and the object appeared to be brighter than on Mar. 7. It could not be detected to a limiting magnitude of R ∼ 17 on Mar. 1. Inspection of spectra (range 310-730 nm) obtained by A. Martel and R. W. Goodrich (University of California at Santa Cruz) on Mar. 9 UT with the Shane 3-m reflector at Lick Observatory reveals that the supernova is of type Ia, probably ∼ 1 week prior to maximum brightness. The minimum of the Si II 635.5-nm trough is at 604 nm, yielding an expansion velocity of roughly 15 000 km/s. The continuum is very blue. This supernova could reach mag 11 to 12, since the parent galaxy is in the Virgo cluster; further observations at all wavelengths are urged.”

COMET McNAUGHT-RUSSELL (1993v) Total visual magnitude estimates from Australia, all made using 10×50

binoculars: Mar. 2.46 UT, 8.6 (G. Garradd, Tamworth, N.S.W.); 3.42, 7.8 (D. A. J. Seargent, The Entrance, N.S.W.); 9.43, 7.8 (R. H. McNaught, Siding Spring Observatory).

1994 March 9 Daniel W. E. Green